OpenStudy (anonymous):

6w^9 * 3w^3 = 18w^12 is that right

OpenStudy (anonymous):

6w9· 3w3 $\mathrm{Multiply\:the\:numbers:}\:6\cdot\:3=18$Multiply the numbers: 6· 3=18 $=18w^9w^3$=18w9w3 $\mathrm{Apply\:exponent\:rule}:\quad\:a^b\cdot\:a^c=a^{b+c}$Apply exponent rule: ab· ac=ab+c $w^9w^3=\:w^{9+3}=\:w^{12}$w9w3= w9+3= w12 $=18w^{12}$=18w12

OpenStudy (anonymous):

so by this u r correct

OpenStudy (anonymous):

yaya

OpenStudy (anonymous):

np

OpenStudy (anonymous):

:*

OpenStudy (anonymous):

<3

OpenStudy (anonymous):

mwah

OpenStudy (anonymous):

mmmmmwwwwwwaaaaahhhhh! @usercode3rror

OpenStudy (anonymous):

ur bae 4 life